OpenStudy (anonymous):

Write an equation of the line that satisfies the given conditions. Give the equation ​(a) in​ slope-intercept form and ​(b) in standard form. m=-3/8 (-4,8)

OpenStudy (dan815):

y-8=-3/8*(x-(-4)) this is a line with slope -3/8 shifted 8 up and 4 left you can rewrite this in slope intercept or standard form

OpenStudy (dan815):

for slope intercept form you must rewrite it in y=mx+b form --------------------------------- y-8=-3/8*(x-(-4)) y=-3/8*(x+4) +8 expand the -3/8 y=-3/8*x + -3/8*4 + 8 y=-3/8x -12/8 + 8 y =-3/8x +52/8

OpenStudy (dan815):

so m = -3/8 b= 52/8 you can put the 52/8 in simpler fractions

OpenStudy (anonymous):

so it would look like: y=-3/8x+1/2

OpenStudy (dan815):

what is 52/8 in simpler terms

OpenStudy (anonymous):

you can't reduce 52/8 because it is a improper fraction and it would be 6.5

OpenStudy (anonymous):



OpenStudy (dan815):

okay then leave it but if u want u can just put it in simpler fractions

OpenStudy (dan815):

since 4 goes into both

OpenStudy (dan815):

52/8= 13/2

OpenStudy (dan815):

now for standard form u have to put it as aX+bY=c a,b,c are constants